OpenStudy (anonymous):

Insert brackets to make these calculations correct. (a) 6 + 2 × 4 = 32 (b) 6 + 2 × 4 − 1 = 12 (c) 6 + 2 × 4^2 = 70

OpenStudy (anonymous):

(6+2) x 4 = 32 6 + (2 x (4-1)) = 12 6 + (2 x 4)^2 = 70